The total profit for a company in Feb. was 20% higher than it was in Jan. The total profit for the two months was $157,498. Find the profit for each month.
I know there is a formula for this was but I can't figure it out.
The answer is Jan. 71,590.00 Feb. 85,908

Answer by checkley75(3666)

LET X BE THE PROFIT FOR JAN. & 1.2X FOR THE PROFIT FOR FEB.
NOW ADD THE 2 MONTHS PROFIT:
X+1.2X=157,498
2.2X=157,498
X=157,498/2.2
X=71,590 IS THE PROFIT FOR JAN.
71,590*1.2=85,908 IS THE PROFIT FOR FEB.
PROOF
71,590+85,908=157,498.